# Global decomposition of networks into multiple cores formed by local hubs

### v1

@@ I Introduction
By discriminating the differential effects of the global and local hubs or cores, it would be possible to significantly enhance our understanding of both structural and dynamical aspects of networks. In this paper, we introduce the local-edge decomposition, which is based on the product of hub centrality values of the nodes connected by the edge of interest. One particularly promising aspect of this hub-centrality-based edge-pruning process is the presence of a natural breaking point between the zero and nonzero edge-importance values, detected by the giant component size, and we take the series of such breaking or cusp points to build a systematic procedure to decompose a network. It assigns nodes’ hierarchical levels and uncovers the onion structure of networks [31].
Because we use the concept of local hubs, our decomposition scheme takes a unique viewpoint of putting local hubs in the highest hierarchical level, and this perspective will open new possibilities of applications. Among the possibilities, we study and propose the core-periphery structure and the score function to find it, based on the nodes’ local-hub-based hierarchical levels. We finalize the paper by extending our method by both zooming-in and zooming-out—core-periphery structures within communities and coarse-graind supernode networks, which may provide a crucial clue to solve the conundrum of the necessity for “something else” in core-periphery [20] by finding multiple cores composed of local hubs.
@@ II.1 Edge pruning and cusp point
To quantify edges’ importance to set the criterion for decomposition, we try edge betweenness centrality related to the shortest-path-based global transport dynamics [34] and the product of nodes’ importance, which are attached to both ends of the edge. In the latter case, the property of an edge that connects nodes $i$ and $j$ is given by the product of values assigned to each edge as
@@
In our edge-pruning strategy, we calculate the set of edge importance $\{\mathscr{E}_{ij}\}$ in the original network and use it throughout the process until the end. We do not recalculate the hub centrality and its product during the edge removal process, as it is not our main interest to investigate the modified structure itself  [^1] Another side effect of such recalculation is that the whole process would depend heavily on the order of removed edges with the same value of $\mathscr{E}_{ij}$, which leads to much more complicated situations in practice that requires ensemble average over multiple realizations.; rather, we would like to extract more central parts in terms of the edge importance in the original structure as we proceed. We implement repeated edge pruning  [^2] We suggest using the heap queue structure in numerical simulations for this., which removes the least important edges at each time step. We first examine a natural measure to characterize the edge-pruning process, which is the relative giant (i.e., the largest connected) component size $G$ of the remaining structure with respect to the original network size, as a function of the fraction $p$ of removed edges. To focus on the fragmentation caused by the edge-pruning process, for all of our numerical studies, we use the giant component of the original networks (the sizes of which are listed inside the parentheses in Table 1) in the beginning.
Table 1: Information of real-world networks considered in the work: number of nodes ($N$), giant component size ($G$), number of edges ($E$), number of edges of the giant component ($E_{G}$), average degree ($\langle k\rangle$), average degree of the giant component ($\langle k\rangle_{G}$), average local clustering coefficient ($c_{l}$), and assortativity ($r$). When implementing local-edge decomposition, it is conducted within the giant component $G$. The collaboration network is the authors’ collaboration network in the field of computational geometry.
Networks $N$ ($G$) $E$ ($E_{G}$) $\langle k\rangle$ ($\langle k\rangle_{G}$) $c_{l}$ $r$ Ref.
@@
A typical example is the case of the collaboration network in the field of computational geometry (CB) [37] (see Table 1 for basic statistics) shown in Fig. 1(a). In the case of $k\mbox{-}\mathrm{P}$, $G$ decreases monotonically. For $C_{B}$, the value of $G$ remains high when $p\lesssim 0.4$, but $G$ decreases rapidly once $p\gtrsim 0.4$. In contrast to previous cases, $h\mbox{-}\mathrm{P}$ shows a cusp point ($p_{c}$), which is the point where the rate of change in $G$ abruptly changes; the slope of $G(p)$ suddenly drops when $p$ crosses the value $p_{c}$. Other networks listed in Table 1 also have the cusp point when the edges are removed by $h\mbox{-}\mathrm{P}$ and show qualitatively similar behavior to Fig. 1(a). This seemingly puzzling behavior is, in fact, simply explained, with the hint from the observation of the fraction $e_{0}$ of edges with exactly null importance, i.e., the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$, which are precisely the edges connecting at least one node with the locally smallest degree: $h_{i}=0$ by the product rule. As shown in Fig. 1(b), for all of the four real networks we examine, the $e_{0}$ value coincides with the cusp point $p_{c}$ for each network. This implies that the transition between $p<p_{c}$ and $p>p_{c}$ corresponds to the part where the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ are removed (by our edge-pruning rule, they must be removed first) versus the part where the rest of the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}>0$ are removed.
@@
Moreover, for any node $i$ with $h_{i}=0$, by definition, all of its neighboring nodes should have larger degrees than $k_{i}$, so $\mathscr{E}^{h\mbox{-}\mathrm{P}}_{ij}=0$ for any node $j$ connected to $i$. As a result, all of the edges connected to $i$ will be removed during the $p<p_{c}$ pruning process. Conversely, as already stated, each of the edge-removal processes for $p<p_{c}$ involves such a node with $h_{i}=0$. Therefore, the edge-pruning process for $p<p_{c}$ exactly corresponds to the process of pinpointing the nodes with $h_{i}=0$ and removing all of their edges. This surgical removal effectively isolates those locally least important nodes and leaves the rest of the network as a new giant component precisely at $p=p_{c}$. In principle, it is possible for a node with nonzero hub centrality to be isolated as a result of the removal of all of its neighbors with $h=0$ even for $p<p_{c}$  [^3] One can imagine a following (extreme) case: assume a node with $h>0$ connected to a number of other nodes, all of which have $h=0$ and are connected to very large and dense clique-like subnetworks on the other side than the original node itself. In that case, by the elimination of the neighboring nodes, the original node will be isolated. This effect is reflected in the actual appearance of $g<1-n_{0}$ in Fig. 2(d), but actually it is quite rare (usually less than $0.1\%$ of nodes in our empirical networks).. However, one can check that the relation $g=1-n_{0}$ holds (under the assumption of the absence of such cases; in general, $g<1-n_{0}$ caused by the possibility of the aforementioned “casualty” node with $h>0$) for our empirical networks, where $n_{0}$ is the fraction of the nodes with $h_{i}=0$ and $g\equiv G(p=p_{c})$ is the fraction of the giant component remaining at $p=p_{c}$, from Fig. 1(c). In other words, the separation of zero-hub-centrality nodes does not cause the noticeable separation of other nodes with nonzero hub centrality from the giant component up to $p=p_{c}$, and most nodes with $h>0$ belong to the new giant component at $p=p_{c}$.
@@ II.2 Local-edge decomposition and node hierarchy
Figure 2: (a)–(c) The giant component size $G$ as a function of the fraction $p$ of removed edges based on the hub-centrality product ($h\mbox{-}\mathrm{P}$). (a) The original collaboration network used in Fig. 1. (b) The backbone of the original network. (c) The backbone of the original network’s backbone. (d) The fraction $n_{0}$ of zero hub-centrality nodes as a function of the relative fraction $g$ of the giant component size with zero hub-centrality nodes removed in decomposed networks. The solid line represents $n_{0}=1-g$, which implies the set of zero hub-centrality nodes $\approx$ the shell at each decomposition level.
@@
To verify our earlier argument $n_{0}\approx 1-g$ (for $p<p_{c}$, the edges are eliminated if and only if $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$) at each level, we show all of the $(n_{0},g)$ pairs calculated at each level on top of the line $n_{0}\approx 1-g$, and one can observe the excellent agreement. The other networks listed in Table 1 also show qualitatively similar results to Fig. 2. The final level is composed of nodes with the same degree and the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ (no node’s degree can be smaller than any other nodes’ degree), so the whole process is naturally terminated by the elimination of the entire edges altogether along with all of the nodes; generally, it forms a clique in real-world networks according to our observation. Therefore, we conclude that a network can be decomposed as an onion-like structure [31], wherein we extract the core of the onion by removing locally least important nodes iteratively through LED.
@@
II.3 Comparison with the k-core decomposition
@@ III.1 Core-periphery structure and score
Basically, it represents an adjacency matrix with the perfect shape. The condition $1$ is quantified with $D_{cc}$, and the condition $2$ is measured by $M_{cp}/N_{p}$. The quantity $M_{cp}$, the number of edges connecting the core and periphery nodes, is also the number of periphery nodes connected to core nodes for a simple network, so the ratio $M_{cp}/N_{p}$ represents the fraction of periphery nodes that are connected to the core nodes. In other words, the extent to which the core and periphery nodes satisfy the above two conditions is determined by $D_{cc}$ and $M_{cp}/N_{p}$, respectively. To quantitatively assess the core-periphery structure using these criteria, we define the core-periphery score as,
@@
where $D=2(M_{cc}+M_{cp}+M_{pp})/[(N_{c}+N_{p})(N_{c}+N_{p}-1)]$ is the density of the entire network. This represents the difference between the actual similarity and the expected similarity by chance as the null-model case, concerning the ideal CP. The core-periphery score is maximized when both conditions are satisfied, thus it helps to identify the most optimal boundary level $L_{B}^{*}$ for distinguishing between the core and periphery of a network. As the boundary level $L_{B}$ increases from $1$ to $18$ ($S_{cp}$ for $L_{B}=0$ is undefined), the distinction between core and periphery nodes becomes clearer, causing the core-periphery score to gradually increase. At $L_{B}^{*}=18$, the core-periphery score reaches its maximum value, and this means that nodes belonging to $L\geq L_{B}^{*}=18$ form the core, while the remaining nodes form the periphery in this network. However, when $L_{B}$ exceeds $L_{B}^{*}=18$, $M_{cp}$ decreases and the core-periphery score starts to decrease.
Despite the existence of optimal boundary $L_{B}^{*}=\mathrm{arg\,max}_{L_{B}}S_{cp}$, the core-periphery score itself from this dichotomous distinction between a single core and the rest as the periphery is quite small ($\lesssim 0.1$) even near $L_{B}^{*}$ in the CB network, as shown in Fig. 6. As previously discussed, it stems from the obvious fact that the assumption of single core-periphery separation has its clear limitation for this large-scale real network. The macroscale network is composed of various mesoscale structural features (multiple hubs, communities, etc.) and their interwoven mixture. Since the LED uses nodes’ local relative-degree information, it simultaneously decomposes the network in various heterogeneous places, e.g., communities with different scales, as discussed in Sec. II.3 with Fig. 5, which causes the low value of $S_{cp}$. Therefore, it is clear that we need to sharpen our method more to properly investigate the organizational structure of a network, and that is the final topic of this paper in the next subsection.

### v2

@@ I Introduction
By discriminating the differential effects of the global and local hubs or cores, it would be possible to significantly enhance our understanding of both structural and dynamical aspects of networks. In this paper, we introduce the local-edge decomposition, which is based on the product of hub centrality values of the nodes connected by the edge of interest. One particularly promising aspect of this hub-centrality-based edge-pruning process is the presence of a natural breaking point between the zero and nonzero edge-importance values, detected by the giant component size, and we take the series of such breaking or cusp points to build a systematic procedure to decompose a network. It assigns nodes’ hierarchical levels and uncovers the onion structure of networks [31], as we demonstrate for real and model networks.
Because we use the concept of local hubs, our decomposition scheme takes a unique viewpoint of putting local hubs at the highest hierarchical level, and this perspective will open new possibilities for applications. Among the possibilities, we study and propose the core-periphery structure and the score function to find it, based on the nodes’ local-hub-based hierarchical levels. We finalize the paper by extending our method by both zooming-in and zooming-out—core-periphery structures within communities and coarse-graind supernode networks, which may provide a crucial clue to solve the conundrum of the necessity for “something else” in core-periphery [20] by finding multiple cores composed of local hubs.
@@ II.1 Edge pruning and cusp point
To quantify edges’ importance to set the criterion for decomposition, we try edge betweenness centrality related to the shortest-path-based global transport dynamics [34] and the product of nodes’ importance, which are attached to both ends of the edge. In the latter case, the property of an edge that connects nodes $i$ and $j$ is given by the product of values assigned to each node as
@@
In our edge-pruning strategy, we calculate the set of edge importance $\{\mathscr{E}_{ij}\}$ in the original network and use it throughout the process until the end. We do not recalculate the hub centrality and its product during the edge removal process, as it is not our main interest to investigate the modified structure itself  [^1] Another side effect of such recalculation is that the whole process would depend heavily on the order of removed edges with the same value of $\mathscr{E}_{ij}$, which leads to much more complicated situations in practice that requires ensemble average over multiple realizations.; rather, we would like to extract more central parts in terms of the edge importance in the original structure as we proceed. We implement repeated edge pruning  [^2] We suggest using the heap queue structure in numerical simulations for this., which removes the least important edges at each time step. We first examine a natural measure to characterize the edge-pruning process, which is the relative size $G=|\mathcal{G}|/N$ of giant (i.e., the largest connected) component $\mathcal{G}$ with respect to the original network size $N$, as a function of the fraction $p$ of removed edges. To focus on the fragmentation caused by the edge-pruning process, for all of our numerical studies, we use the giant component of the original networks (the sizes of which are listed inside the parentheses in Table 1) in the beginning.
Table 1: Information of real-world networks considered in the work: number of nodes ($N$), the size ($|\mathcal{G}|$) of the giant component $\mathcal{G}$, number of edges ($E$), number of edges of the giant component ($E_{G}$), average degree ($\langle k\rangle$), average degree of the giant component ($\langle k\rangle_{G}$), average local clustering coefficient ($c_{l}$), and assortativity ($r$). When implementing local-edge decomposition, it is conducted within the giant component $G$. The collaboration network is the authors’ collaboration network in the field of computational geometry.
Networks $N$ ($|\mathcal{G}|$) $E$ ($E_{G}$) $\langle k\rangle$ ($\langle k\rangle_{G}$) $c_{l}$ $r$ Ref.
@@
A typical example is the case of the collaboration network in the field of computational geometry (CB) [37] (see Table 1 for basic statistics) shown in Fig. 1(a). In the case of $k\mbox{-}\mathrm{P}$, $G$ decreases monotonically. For $C_{B}$, the value of $G$ remains high when $p\lesssim 0.4$, but $G$ decreases rapidly once $p\gtrsim 0.4$. Notably, there is a cusp-like point at $p\approx 0.8$, which is the point where the rate of change in $G(p)$ qualitatively changes. For $h\mbox{-}\mathrm{P}$, it shows a more prominent cusp point at $p=p_{c}\approx 0.5$, and the slope of $G(p)$ suddenly drops almost to zero when $p$ crosses the value $p_{c}$. As we will discuss later, in the case of $h\mbox{-}\mathrm{P}$, a significant structural change in the network occurs at the cusp point, whereas we do not observe such a significant change for $C_{B}$. Other networks listed in Table 1 also have the clear-cut cusp point when the edges are removed by $h\mbox{-}\mathrm{P}$ and show qualitatively similar behavior to Fig. 1(a). This seemingly puzzling behavior is, in fact, simply explained, with the hint from the observation of the fraction $e_{0}$ of edges with exactly null importance, i.e., the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$, which are precisely the edges connecting at least one node with the locally smallest degree: $h_{i}=0$ by the product rule. As shown in Fig. 1(b), for all of the four real networks we examine, the $e_{0}$ value coincides with the cusp point $p_{c}$ for each network. This implies that the transition between $p<p_{c}$ and $p>p_{c}$ corresponds to the part where the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ are removed (by our edge-pruning rule, they must be removed first) versus the part where the rest of the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}>0$ are removed.
@@
Moreover, for any node $i$ with $h_{i}=0$, by definition, all of its neighboring nodes should have larger degrees than $k_{i}$, so $\mathscr{E}^{h\mbox{-}\mathrm{P}}_{ij}=0$ for any node $j$ connected to $i$. As a result, all of the edges connected to $i$ will be removed during the $p<p_{c}$ pruning process. Conversely, as already stated, each of the edge-removal processes for $p<p_{c}$ involves such a node with $h_{i}=0$. Therefore, the edge-pruning process for $p<p_{c}$ exactly corresponds to the process of pinpointing the nodes with $h_{i}=0$ and removing all of their edges. This surgical removal effectively isolates those locally least important nodes and leaves the rest of the network as a new giant component precisely at $p=p_{c}$. In principle, it is possible for a node with nonzero hub centrality to be isolated as a result of the removal of all of its neighbors with $h=0$ even for $p<p_{c}$  [^3] One can imagine a following (extreme) case: assume a node with $h>0$ connected to a number of other nodes, all of which have $h=0$ and are connected to very large and dense clique-like subnetworks on the other side than the original node itself. In that case, by the elimination of the neighboring nodes, the original node will be isolated. This effect is reflected in the actual appearance of $g\equiv G(p=p_{c})<1-n_{0}$ in Fig. 2(d), but actually it is quite rare (usually less than $0.1\%$ of nodes in our empirical networks).. However, one can check that the relation $g\equiv G(p=p_{c})=1-n_{0}$ holds (under the assumption of the absence of such cases; in general, $g<1-n_{0}$ caused by the possibility of the aforementioned “casualty” node with $h>0$) for our empirical networks, where $n_{0}$ is the fraction of the nodes with $h_{i}=0$ and $g$ is the fraction of the giant component remaining at $p=p_{c}$, from Fig. 1(c). In other words, the separation of zero-hub-centrality nodes does not cause the noticeable separation of other nodes with nonzero hub centrality from the giant component up to $p=p_{c}$, and most nodes with $h>0$ belong to the new giant component at $p=p_{c}$.
@@ II.2 Local-edge decomposition and node hierarchy
Figure 2: (a)–(c) The giant component size $G$ as a function of the fraction $p$ of removed edges based on the hub-centrality product ($h\mbox{-}\mathrm{P}$). (a) The original collaboration network used in Fig. 1. (b) The backbone of the original network. (c) The backbone of the original network’s backbone. (d) The relative fraction $g$ of the giant component size with zero hub-centrality nodes removed as a function of the fraction $n_{0}$ of zero hub-centrality nodes in decomposed networks. The solid line represents $n_{0}=1-g$, which implies the set of zero hub-centrality nodes $\approx$ the shell at each decomposition level.
@@
Figure 2(d) demonstrates that our earlier argument holds well; $n_{0}\approx 1-g$ (for $p<p_{c}$) and the edges are eliminated if and only if $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ at each level. As a result, all of the $(n_{0},g)$ pairs at different levels fall on top of the $n_{0}=1-g$ line. The other networks listed in Table 1 also show qualitatively similar results to Fig. 2. The final level is composed of nodes with the same degree and the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ (no node’s degree can be smaller than any other nodes’ degree), so the whole process is naturally terminated by the elimination of the entire edges altogether along with all of the nodes; generally, it forms a clique in real-world networks according to our observation. Therefore, we conclude that a network can be decomposed as an onion-like structure [31], wherein we extract the core of the onion by removing locally least important nodes iteratively through LED.
@@
II.3 Local-edge decomposition of the Barabási-Albert model
Figure 4: (a) The giant component size $G$ as a function of the fraction $p$ of removed edges based on $h\mbox{-}\mathrm{P}$, applied to a single realization of the BA model [44] with $10^{5}$ nodes and the number $m=4$ of stubs for each newly added node. The black squares indicate the original network. The red circles, blue triangles, and purple inverted triangles represent the results obtained for the primary, secondary, and tertiary backbones, respectively, of the original network. (b) The relative fraction $g$ of the giant component size with zero hub-centrality nodes removed as a function of the fraction $n_{0}$ of zero hub-centrality nodes in decomposed networks at different levels. The black squares, red circles, and blue triangles represent the cases with $m=2$, $4$, and $6$, respectively. We only plot the cases where the backbone with the number of nodes $>1\%$ of that of the original network.
Our LED analysis of real networks may give the impression that it is not applicable to simple model networks without mesoscale heterogeneity. This assumption must be correct for a completely homogeneous mixture of random connections represented by the Erdős-Rényi (ER) random graph [45] as all of the nodes there are topologically equivalent. In this subsection, however, we provide an example of a model network where its construction principle induces the sequential hierarchy of LED levels that is well-detected by our method. The historically important and celebrated Barabási-Albert (BA) model [44] is well-known for its power-law degree distribution that gave birth to the famous notion of “scale-free” networks (SFNs). At the same time, however, it is relatively less emphasized that one of the key factors of the model, the growth by attaching nodes with a fixed number $m$ of stubs, provides the intrinsic correlation between individual nodes’ time of inception and its degree [7, 46], statistically speaking. As a result, in contrast to other SFN models without the “growing” mechanism [47, 48], the BA model tends to be organized as in Fig. 3, where central (peripheral) nodes correspond to older (newer) nodes.
Such an organizational structure is well-captured by the hub-centrality-based LED applied to the BA model, as shown in Fig. 4, where we show the $G(p)$ and $g(n_{0})$ curves for different LED levels as in the cases of real networks. The evidence cusp points in Fig. 4(a) and the relation $g\approx 1-n_{0}$ at each cusp point in Fig. 4(b) indicate that our method successfully reveals the aforementioned organizational structure of the model. The most outstanding feature about the BA-model case is the fact that $G(p)$ curves for different levels are collapsed onto a single characteristic curve. This implies that in contrast to real networks [see Fig. 2(a)–(c)] each LED level or backbone of the BA model has almost the same organizational structure starting from the original network, which is also reflected in the collapsed points $\left(n_{0},g(n_{0})\approx 1-n_{0}\right)$ for different levels at each $m$ value [Fig. 4(b)]. This self-similar or scale-invariant property reminds us of more fundamental scale-free property than just the power-law degree distribution [49]. In retrospect, our edge-pruning processes correspond to the time-reversal processes of the node and edge growing mechanism imposed in the creation of the BA model on average (the preferential attachment for creation versus the preferential detachment [50] for LED), so the result is not too surprising. Moreover, even at a single level, in addition to the primary cusp point at $p\approx 0.4$, there are (less apparent) secondary ($p\approx 0.6$) and more cusp-like points in the BA model (again, at almost exactly the same $p$ for every level). The first cusp point is the point that the nodes with zero hub centrality are separated from the giant component as in real networks, but the others are causes by the collective and systematic decrement of the nodes with low hub-centrality values, which are absent in real networks without such an artificial systematic attaching mechanism of nodes with a fixed number of stubs.
II.4 Comparison with the k-core decomposition
@@ III.1 Core-periphery structure and score
Basically, it represents an adjacency matrix with the perfect shape. The condition $1$ is quantified with $D_{cc}$, and the condition $2$ is measured by $M_{cp}/N_{p}$. The quantity $M_{cp}$, the number of edges connecting the core and periphery nodes, is also the number of periphery nodes connected to core nodes for a simple network without the possibility of multiple edges between a pair of nodes, so the ratio $M_{cp}/N_{p}$ represents the fraction of periphery nodes that are connected to the core nodes. In other words, the extent to which the core and periphery nodes satisfy the above two conditions is determined by $D_{cc}$ and $M_{cp}/N_{p}$, respectively. To quantitatively assess the core-periphery structure using these criteria, we define the core-periphery score as,
@@
where $D=2(M_{cc}+M_{cp}+M_{pp})/[(N_{c}+N_{p})(N_{c}+N_{p}-1)]$ is the density of the entire network. This represents the difference between the actual similarity and the expected similarity by chance as the null-model case, concerning the ideal CP. The core-periphery score is maximized when both conditions are satisfied, thus it helps to identify the most optimal boundary level $L_{B}^{*}$ for distinguishing between the core and periphery of a network. When a network exhibits an ideal CP structure, $S_{cp}$ is close to $1$. In contrast, in a null model network, $S_{cp}=0$ because the core-periphery score represents the comparison of the network with the null model by definition. As the boundary level $L_{B}$ increases from $1$ to $18$ ($S_{cp}$ for $L_{B}=0$ is undefined), the distinction between core and periphery nodes becomes clearer, causing the core-periphery score to gradually increase. At $L_{B}^{*}=18$, the core-periphery score reaches its maximum value, and this means that nodes belonging to $L\geq L_{B}^{*}=18$ form the core, while the remaining nodes form the periphery in this network. However, when $L_{B}$ exceeds $L_{B}^{*}=18$, $M_{cp}$ decreases and the core-periphery score starts to decrease.
Despite the existence of optimal boundary $L_{B}^{*}=\mathrm{arg\,max}_{L_{B}}S_{cp}$, the core-periphery score itself from this dichotomous distinction between a single core and the rest as the periphery is quite small ($\lesssim 0.1$) even near $L_{B}^{*}$ in the CB network, as shown in Fig. 7. As previously discussed, it stems from the obvious fact that the assumption of single core-periphery separation has its clear limitation for this large-scale real network. The macroscale network is composed of various mesoscale structural features (multiple hubs, communities, etc.) and their interwoven mixture. Since the LED uses nodes’ local relative-degree information, it simultaneously decomposes the network in various heterogeneous places, e.g., communities with different scales, as discussed in Sec. II.4 with Fig. 6, which causes the low value of $S_{cp}$; both $D_{cc}$ and $N_{cp}/N_{p}$ cannot have high values simultaneously.